Jenny is planning to deposit 17,000.00 pesos Quezon Metropolitan Bank is offering 7.5% compounded semi-annually for 5 years while Quezon Premier Bank is offering 7% compounded monthly for 5 years. Wich bank should she deposit her money? Justify you answer by computing and comparing the maturity value for each bank.


1
Expert's answer
2021-11-29T01:48:50-0500

Compound interest


"A=P(1+\dfrac{r}{n})^{nt}"

Given "P=17000, t=5"


Quezon Metropolitan Bank: "r=0.075, n=2"


"A=17000(1+\dfrac{0.075}{2})^{2(5)}=24565.75"

Quezon Premier Bank: "r=0.07, n=12"


"A=17000(1+\dfrac{0.07}{12})^{12(5)}=24099.63"

She should deposit her money Quezon Metropolitan Bank .